OpenStudy (anonymous):

convert from degrees to radians. Use the value of π found on a calculator and round answers to four decimal places ,as needed. 36°

OpenStudy (anonymous):

The answer is π/5. I want to know how this is the answer.

OpenStudy (anonymous):

We need to take a number from the range of \[[0,2\pi]\] and map it, in a linear manner to\[[0,360]\]So, we get\[360k=2\pi\]Where k is our conversion factor, solving for k\[k=\frac{\pi}{180}\]So, if we have 36 degrees, we convert to radians accordingly by multiplying by k. This means\[36k=36\cdot\frac{\pi}{180}=\frac{\pi}{5}\]